Automotive Products (AP) designs and produces automotive parts. In 2017, actual manufacturing overhead is $317,250. APs simp-X i Requirements 1. 2. 3. Compute the manufacturing overhead allocated to each customer in 2017 using the simple costing sysX Х Data Table Usage of Cost Drivers by Customer Contract Northern Hamlet Motors Motors Jarrin Auto Manufacturing Overhead inRequirement 1. Compute the manufacturing overhead allocated to each customer in 2017 using the simple costing system that useRequirement 1. Compute the manufacturing overhead allocated to each customer in 2017 using the simple costing system that useRequirement 2. Compute the manufacturing overhead allocated to each customer in 2017 using department-based manufacturing oveRequirement 2. Compute the manufacturing overhead allocated to each customer in 2017 using department-based manufacturing oveAPs 2017 manufacturing overhead rates by department are as follows: V Manufacturing Overhead Rate Department (by Department)Requirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 3. Comment on your answers in requirements 1 and 2. Which customer do you think was complaining about being overcRequirement 4. How else might AP use the information available from its department-by-department analysis of manufacturing ov
Answer #1
1 Manufacturing overhead rate=Total manufacturing overhead/Total machine hours=317250/(90+2000+910)=317250/3000=$105.75 per machine hour
Overhead allocated=Machine hours used*Manufacturing overhead rate
Customer Machine hours used Overhead allocated
a a*105.75
Northern motors 90 9518
Hamlet motors 2000 211500
Jarrin auto 910 96233
2 Manufacturing overhead by department=Manufacturing overhead in the department/Total cost drivers
Department Manufacturing overhead Cost driver Total cost driver Manufacturing overhead rate by department
a b a/b
Design 44100 CAD-Design-hours 420 105 Per CAD-Design-hour
(120+210+90)
Engineering 33150 Engineering-hours 390 85 Per Engineering-hour
(80+50+260)
Production 240000 Machine-hours 3000 80 Per machine-hour
Department Northern motors Hamlet motors Jarrin auto Total
Manufacturing overhead rate Usage of cost driver Overhead allocated Manufacturing overhead rate Usage of cost driver Overhead allocated Manufacturing overhead rate Usage of cost driver Overhead allocated
a b c=a*b d e f=d*e g h i=g*h c+f+i
Design 105 120 12600 105 210 22050 105 90 9450 44100
Engineering 85 80 6800 85 50 4250 85 260 22100 33150
Production 80 90 7200 80 2000 160000 80 910 72800 240000
26600 186300 104350 317250
3 Overhead allocated:
Northern motors Hamlet motors Jarrin auto
Simple-costing system 9518 211500 96233
Department-based rates 26600 186300 104350
Under the simple-costing system,Hamlet motors was likely unhappy about their bill,because the contract appears to have been overpriced.
If the new department-based rates are used to price contracts,Hamlet motors will likely be happier about their bill
If the new department-based rates are used to price contracts,Northern motors will likely be unhappy.
This is because their bills under the simple-costing system appear to have been underpriced.
Using the department-based rates,their bills will be overpriced.
If the new department-based rates are used,AP should explain to the unhappy customers how the calculation was done
and point out their high use of cost drivers relative to other customers
4 B. to reevaluate employee pay scales
C. to examine and streamline its own operations so that there is maximum value- added from all indirect resources
D. to set targets over time to reduce both the consumption of each indirect resource and the unit costs of the resources
5 D. All of the above